# An abstract framework for heterogeneous coupling: stability, approximation and applications

### v1

@@ 4. Solving the resulting discrete problem
$\begin{array}[]{lcl}\sigma(\eta,\mu)+b(\widetilde{z},\mu)&=&\sigma(\lambda,\mu),\qquad\forall\mu\in\Lambda_{\delta}\\ b(z^{\prime},\eta)&=&0,\qquad\forall z^{\prime}\in Z.\end{array}$
Reduction to $\widehat{\Lambda}_{\delta}$
Letting $\lambda$ be the solution to problem 3.1, we have that $\lambda^{0}=\lambda-\Pi_{\sigma}\lambda$ is the first component of the solution $(\lambda^{0},z^{0})\in\Lambda_{\delta}\times Z$ to
(4.1) $\begin{array}[]{lcl}\sigma(\lambda^{0},\mu)-b(z^{0},\mu)&=&0,\qquad\forall\mu\in\Lambda_{\delta}\\ b(z^{\prime},\lambda^{0})&=&-\langle f,z^{\prime}\rangle,\qquad\forall z^{\prime}\in Z.\end{array}$
We assume that the scalar product $\sigma$ is chosen in such a way that such a problem can be solved cheaply (at least with respect to the cost of the solution of the whole Problem 3.1), and $\widehat{\lambda}=\Pi_{\sigma}\lambda=\lambda-\lambda^{0}$ is the solution to the following reduced problem
We know by Lemma 3.6 that $s_{h}$ is coercive on $\widehat{\Lambda}_{\delta}$, and, consequently, that the corresponding operator $\widehat{S}_{h}:\widehat{\Lambda}_{\delta}\to\widehat{\Lambda}_{\delta}^{\prime}$ is invertible. The problem can be then addressed by any iterative solver well suited to solve linear systems with positive definite matrices. For instance, if, in addition, the operator $A^{+}_{h}$ is symmetric, which implies that $S_{h}$ and, consequently, $\widehat{S}_{h}$ are symmetric, one can resort to the preconditioned conjugate gradient method.
Construction of the preconditioner $\widehat{M}$
What we wold like, in order to build a preconditioner, is something resembling an inverse of $S_{h}$, if $S_{h}$ was invertible. If $A^{+}_{h}$ and $B_{\delta}$ were also, in turn, invertible, this would be
(4.3) $(BA^{+}_{h}B^{T})^{-1}=B^{-T}(A^{+}_{h})^{-1}B^{-1}.$
Of course, none of the above mentioned operators is generally invertible; the idea is then to use the right-hand side expression as a guideline, and replace $B^{-1}$ and $(A_{h}^{+})^{-1}$, respectively, with a suitable pseudo inverse $B_{\delta}^{+}$, and with $A$ (or some spectrally equivalent $\widetilde{A}$).
Following the approach put forward by the FETI preconditioner, we rely on three ingredients The first ingredient is a finite dimensional space $V_{h}\subseteq V$, which we assume to satisfy
(4.4) $Z\subseteq V_{h},\qquad\text{and}\qquad\inf_{\lambda\in\Lambda_{\delta}}\sup_{v\in V_{h}}\frac{b(v,\lambda)}{\|v\|_{V}\|\lambda\|_{\Lambda}}\gtrsim 1.$
Observe that, if a space $C_{h}$ underlying the approximate pseudo inverse $A^{+}_{h}$ is available, a natural choice for $V_{h}$ is $V_{h}=C_{h}$, but we are also interested in situations in which such a space might not be accessible. Also note that $V_{h}$ needs not satisfy any approximation assumption.
We let $A_{\delta}:V_{h}\to V_{h}^{\prime}$ and $B_{\delta}:V_{h}\to\Lambda_{\delta}^{\prime}$, be the discrete versions of the operators $A$ and $B$, defined as
$\langle A_{\delta}v,w\rangle=a(v,w)\ \forall v,w\in V_{h},\qquad\langle B_{\delta}v,\lambda\rangle=b(v,\lambda)\ \forall v\in V_{h},\lambda\in\Lambda_{\delta}.$
The second ingredient is a discrete scalar product $d:V_{h}\times V_{h}\to\mathbb{R}$ (not necessarily equivalent to the natural scalar product for $V$) and a corresponding $d$-orthogonal splitting of the space $V_{h}$ as
$V_{h}=\ker B_{\delta}\oplus(\ker B_{\delta})^{\perp},$
where
$\ker B_{\delta}=\{v\in V_{h}:\ b(v,\mu)=0,\ \forall\mu\in\Lambda_{\delta}\},$
$(\ker B_{\delta})^{\perp}=\{v\in V_{h}:\ d(v,w)=0,\ \forall w\in\ker B_{\delta}\}.$
We let $\Pi_{d}:V_{h}\to(\ker B_{\delta})^{\perp}$ denote the $d$-orthogonal projection onto $(\ker B_{\delta})^{\perp}$. Whichever the choice of the scalar product $d$, the restriction of $B_{\delta}$ to $(\ker B_{\delta})^{\perp}$ is bijective, and we let $B_{\delta}^{+}:\Lambda_{\delta}^{\prime}\to(\ker B_{\delta})^{\perp}$ denote its inverse, which is a pseudo inverse of $B_{\delta}$, verifying
We easily realize that, given $\phi\in\Lambda_{\delta}^{\prime}$, the pseudo inverse $B_{\delta}^{+}\phi$ can be defined as $B_{\delta}^{+}\phi=v$, where $(v,\zeta)\in V_{h}\times\Lambda_{\delta}$ is the solution to
$\begin{array}[]{lcl}d(v,w)+b(w,\zeta)&=&0,\qquad\forall w\in V_{h}\\ b(v,\mu)&=&\langle\phi,\mu\rangle,\forall\mu\in\Lambda_{\delta}.\end{array}$
Following the guidelines provided by the formal inversion expression (4.3) we introduce the linear operator $M:\Lambda_{\delta}^{\prime}\to\Lambda_{\delta}$, defined as
$M=(B_{\delta}^{+})^{T}A_{\delta}B_{\delta}^{+},$
which corresponds to the bilinear form $m:\Lambda_{\delta}^{\prime}\times\Lambda_{\delta}^{\prime}$ defined as
$m(\phi,\psi)=a(B_{\delta}^{+}\phi,B_{\delta}^{+}\psi).$
We finally define our preconditioner $\widehat{M}:\widehat{\Lambda}_{\delta}^{\prime}\to\widehat{\Lambda}_{\delta}$ as the operator corresponding to the bilinear form $\widehat{m}:\widehat{\Lambda}_{\delta}^{\prime}\times\widehat{\Lambda}_{\delta}^{\prime}\to\mathbb{R}$, defined as
$\widehat{m}({\widetilde{\varphi}},\widetilde{\psi})=m(\mathcal{K}{\widetilde{\varphi}},\mathcal{K}\widetilde{\psi}).$
We have the following theorem, the proof of which is fundamentally the same as for the FETI method in domain decomposition framework (see [29]). For the sake of completeness we report it in Appendix B.
Theorem 4.2.
Let $\Pi_{d}:V_{h}\to V_{h}$ denote the ${d}$-orthogonal projection operator onto the ${d}$-orthogonal complement $(\ker B_{\delta})^{\perp}$ of $\ker B_{\delta}$ in $V_{h}$. Then it holds
${\kappa(\widehat{M}\widehat{S}_{h})\lesssim C(V_{h},d,\sigma)}$

### v2

@@ 4.3. The solution recipe
$M^{+}=(B_{\delta}^{+})^{T}\widetilde{A}_{\delta}B_{\delta}^{+}=\sum_{\ell}\sum_{\ell^{\prime}\in\mathcal{N}(\ell)}I^{\Lambda}_{\ell^{\prime}}D^{-T}_{\ell^{\prime}}\underbrace{\left(\begin{array}[]{c}R^{V}_{k^{+}_{\ell^{\prime}}}\\ R^{V}_{k^{-}_{\ell^{\prime}}}\end{array}\right)\widetilde{A}_{\delta}\left(\begin{array}[]{cc}I^{V}_{{k^{+}_{\ell}}}&I^{V}_{{k^{-}_{\ell}}}\end{array}\right)}_{I}D^{-1}_{\ell}R^{\Lambda}_{\ell}$
Remark that for $k\not=k^{\prime}$ we have that $R^{V}_{k^{\prime}}\widetilde{A}_{\delta}I^{V}_{k}=0$, while $R^{V}_{k}\widetilde{A}_{\delta}I^{V}_{k}=\widetilde{A}_{\delta}^{k}$, which eventually simplifies $I$ for $\ell\not$$\ell^{\prime}$. The operator $M^{+}$ is then combined with $\mathcal{K}$ to form the preconditioner $\widehat{M}^{-1}=\mathcal{K}^{T}M^{+}\mathcal{K}$.
5. Examples
The above framework has multiple possible applications, and it can be used for the design of coupling approaches in multiphysics, multiscale or mixed-dimensional problems. The coupling of FEM and BEM [10], the Arlequin method ([25]), the surface/bulk coupling arising in proton transport across biological membranes [46] or in the study of the electrostatic potential in one-layer material semiconductor devices [33], the flux in branching fracture networks [30, 8], can all be cast within our abstract framework, which allows to use previously implemented and optimized solvers of possibly heterogeneous nature for the solution of the individual problems. Of course, the most obvious application is the one to domain decomposition, from which we openly drew inspiration in designing our coupling approach. In the following sections we will show how casting two known domain decomposition approaches in the abstract framework allows first and foremost to design algorithms aimed at coupling black-box subdomain solvers, but also to design new variants of some known preconditioner.
5.1. Domain decomposition with Neumann coupling
Let $\Omega\subset\mathbb{R}^{n}$, $n=2,3$ be a polygonal/polyhedral domain with boundary $\partial\Omega$. We consider the following simple model problem: given $f\in L^{2}(\Omega)$, find $u$ satisfying
[…]
We let $\Omega$ be split as the non-overlapping union of polygonal/polyhedral subdomains: $\overline{\Omega}=\cup_{k=1}^{K}\overline{\Omega}^{k}$ and set $\Gamma^{k}=\partial\Omega^{k}$, $\Gamma=\cup\Gamma^{k}$ denoting the skeleton of the decomposition. We assume that the subdomain $\Omega^{k}$ are chosen in such a way that $\kappa|_{\Omega^{k}}=\kappa_{k}$, $\kappa_{k}\in\mathbb{R}$ positive constants, possibly very different from each other. For each $k$ we let
$V_{k}=\{u\in H^{1}(\Omega^{k}),~u|_{\partial\Omega\cap\Gamma^{k}}=0\},\qquad\|u\|_{V^{k}}=\kappa_{k}^{1/2}\|u\|_{1,\Omega^{k}}.$
Letting $\nu^{k}$ denote the outer unit normal to the subdomain $\Omega^{k}$, on the skeleton $\Gamma$ we next select a normal direction $\nu$, taking care that on $\partial\Omega$ $\nu$ coincides with the outer normal to $\Omega$. We let $\Lambda$ denote the closure of $L^{2}(\Gamma)$ with respect to the norm
$\|\lambda\|_{-1/2,\Gamma}^{2}=\sum_{k}\kappa_{k}^{-1}\|(\nu\cdot\nu^{k})\lambda\|_{-1/2,\Gamma_{k}}^{2},\qquad\|\lambda^{k}\|_{-1/2,\Gamma_{k}}=\sup_{v\in V_{k}}\frac{\int_{\Gamma_{k}}\lambda^{k}v}{\|v\|_{1,\Omega^{k}}}.$
The space $\Lambda$ is well suited to describe the trace on $\Gamma$ of the normal flux $\lambda=\kappa\nabla u\cdot\nu$ of the solution to our problem (we recall that while $\kappa$ and $\nabla u$ may jump across $\Gamma$, $\kappa\nabla u\cdot\nu$ is continuous, so that $\lambda$ is well defined). Equation (5.1) can be split as a system of coupled local problems as: find $(u^{k})_{k}\in V=\prod_{k}V_{k}$, $\lambda\in\Lambda$, such that
(5.2) $\int_{\Omega^{k}}\kappa_{k}\nabla u^{k}\cdot\nabla v^{k}-\int_{\Gamma^{k}}(\nu\cdot\nu^{k})\lambda v^{k}=\int_{\Omega^{k}}fv^{k},\ \text{ for all }v^{k}\in V_{k},\ k\in\{1,\ldots,K\},$
(5.3) $\sum_{k}\int_{\Gamma^{k}}u^{k}(\nu\cdot\nu^{k})\mu=0,\ \text{ for all }\mu\in\Lambda.$
We can now approximate $\Lambda$ with discontinuous piecewise polynomials of order $l$ on a shape regular mesh $\mathcal{K}_{\Gamma}$. In the subdomains we consider arbitrary discretizations yielding approximate solvers $A_{h,k}^{+}$, which we will treat as black boxes, only assuming that Assumption 3.3 holds. These discretizations need neither be of the finite element family, nor they need be of the same type in all subdomains. We could consider meshless methods or even linear neural network discretizations, such as extreme learning machines, in some subdomains and finite elements in other subdomains. Lemma 3.6 implies that, provided $h/\delta<C_{\Lambda}$, Problem 3.1 is well posed and it can be solved by the approach proposed in Section 4. Efficiency will be attained by constructing a preconditioner according the recipe presented therein.
Remark 5.1.
We could also use continuous polynomials, only allowing discontinuities at the cross points (in two dimensions) or on the wirebasket (in three dimensions), where we also modify the space $\Lambda_{\delta}$ in the spirit of the mortar method [7, 4].
We build the bilinear form $\sigma^{*}$ according to the strategy suggested in Remark 4.3. The space $\Lambda_{0}$ can be chosen as the space of discontinuous functions assuming constant values on the edges (in 2D) or faces (in 3D) of the subdomains. A bilinear form $\sigma$ satisfying (4.5) can be defined as follows:
(5.4) $\sigma(\lambda,\mu)=\sum_{k}\kappa_{k}^{-1}\left(H_{k}^{-n}\left(\int_{\partial\Omega^{k}}(\nu\cdot\nu^{k})\lambda\right)\left(\int_{\partial\Omega^{k}}(\nu\cdot\nu^{k})\mu\right)+H\int_{\partial\Omega^{k}}\lambda\mu\right).$
The following lemma, where the implicit constants depend on the shape regularity of the subdomains, but not on their number or size, can be proven by exploiting standard inverse inequalities and [14, Lemma 2.1].
Lemma 5.2.
For all $\lambda\in\Lambda_{0}$ it holds that
$\sigma(\lambda,\lambda)\simeq\|\lambda\|_{\Lambda}^{2}.$
The first step in constructing the preconditioner is choosing the auxiliary spaces $V_{\delta}^{k}$ and the bilinear forms $\widetilde{a}^{k}$. The most natural choice is to introduce the mesh $\mathcal{K}^{k}$ on $\partial\Omega^{k}$ induced by the mesh $\mathcal{K}_{\Gamma}$, and a mesh $\mathcal{T}_{\delta}^{k}$ on $\Omega^{k}$, whose trace on $\Gamma_{k}$ coincides with $\mathcal{K}^{k}$. We then choose $V_{\delta}^{k}$ as the set of order $l+n$ finite elements on $\mathcal{T}_{\delta}^{k}$, and we can set $\widetilde{a}^{k}=a^{k}$. It is not difficult to realize that this choice leads to what is essentially a FETI preconditioner with “deluxe” scaling, the main difference being the fact that, as in a black box framework, the spaces underlying the local black box solvers $A_{h,k}^{+}$ might not be available, rather than relying on such spaces to construct the preconditioners, we rely on the spaces $V_{\delta}^{k}$. Remark that the evaluation of the scalar product $d^{k}$ defined by (4.15) requires solving local problems in the space $V_{0}^{k}$ of “interior” functions, for the construction of the spaces $V^{k}_{\ell}$. Remark also that the mesh $\mathcal{T}_{\delta}^{k}$ need not be quasi uniform and it can be taken much coarser in the interior of $\Omega^{k}$, thus making the preconditioner cheaper.
We can also consider an alternative, inspired by a combination of the theories of Virtual Elements and of substructuring preconditioners à la Bramble Pasciak Schatz. Depending on whether we are in two or three dimensions, we let $V_{\delta}^{k}$ be defined as follows:
  1. a)
    for $d=2$ we set
    (5.5) $V_{\delta}^{k}=\{v\in H^{1}(\Omega):v|_{e}\in\mathbb{P}_{l+2}(e)\,\forall e\in\mathcal{K}^{k},\Delta v=0\ \text{in }\Omega^{k}\};$
  2. b)
    for $d=3$, letting $\mathcal{E}_{\delta}^{k}$ denote the set of edges of the boundary mesh $\mathcal{K}_{\Gamma}$, we set
    (5.6) $V_{\delta}^{k}=\{v\in H^{1}(\Omega):v|_{e}\in\mathbb{P}_{1}(e)\ \forall e\in\mathcal{E}_{\delta}^{k},\\ v|_{f}\in\mathbb{P}_{l+3}(f)\ \forall f\in\mathcal{K}^{k},\ \Delta v=0\ \text{in }\Omega^{k}\}.$
We now let $\widetilde{a}^{k}$ be defined as
(5.7) $\widetilde{a}^{k}(v,w)=\kappa_{k}\sum_{\ell\in\partial(k)}\left(\widetilde{a}^{k}_{\ell}(v,w)+\left(\fint_{\Gamma^{k}_{\ell}}v-\fint_{\Gamma^{k}}v\right)\left(\fint_{\Gamma^{k}_{\ell}}w-\fint_{\Gamma^{k}}w\right)\right),$
where $\Gamma^{k}_{\ell}$, $\ell\in\partial(k)$ are the edges (in 2D) or faces (in 3D) of the subdomain $\Omega^{k}$, and where
(5.8) $\widetilde{a}^{k}_{\ell}(v,v)\simeq|v|^{2}_{1/2,\Gamma^{k}_{\ell}}.$
The construction of bilinear forms $\widetilde{a}^{k}_{\ell}$ satisfying (5.8) is out of the scope of this paper and we refer to [24, 16] for different options stemming from the theory of virtual elements and of substructuring preconditioners. We have the following Lemma.
Lemma 5.3.
For all $v\in V_{\delta}^{k}$ it holds that
$\log(H/\delta)^{-2}\kappa_{k}|v|_{1,\Omega^{k}}^{2}\lesssim\widetilde{a}^{k}(v,v)\lesssim\kappa_{k}|v|_{1,\Omega^{k}}^{2}.$
This has been proven in two dimensions in [11], but the proof there holds unchanged also in three dimensions. In order to define the pseudo inverse $B_{\delta}^{+}$ we need to identify the subspaces $V^{k}_{\ell}$, $\ell\in\partial(k)$. To this aim let $\widetilde{V}^{k}_{\ell}=\{v\in V^{k}_{\delta}:v=0\ \text{ on }\Gamma^{k}\setminus\Gamma^{k}_{\ell}\}$ and let $\pi_{\ell}:V^{k}\to\widetilde{V}^{k}_{\ell}$ be defined as the first component of the couple $(\pi_{\ell}\phi,\lambda)\in\widetilde{V}^{k}_{\ell}\times\Lambda^{m}$ such that for all $(\psi,\mu)\in\widetilde{V}^{k}_{\ell}\times\Lambda^{m}$
$\kappa_{k}\widetilde{a}^{k}_{\ell}(\pi_{\ell}v,\psi)-\int_{\Gamma^{k}_{\ell}}\lambda\psi=0,\qquad\int_{\Gamma^{k}_{\ell}}\pi_{\ell}\phi\mu=\int_{\Gamma^{k}_{\ell}}\phi\mu.$
As we have that
$\inf_{\lambda\in\Lambda^{m}}\sup_{\phi\in\widetilde{V}^{k}_{\ell}}\frac{\int_{\Gamma^{k}_{\ell}}\lambda\phi}{\|\lambda\|_{-1/2,\Gamma^{k}_{\ell}}\|\phi\|_{H^{1/2}_{00}(\Gamma^{k}_{\ell})}}\gtrsim 1,$
$\pi_{\ell}$ is well defined. The following lemma holds, where we let $\Im(X)$ denote the image of the operator $X$.
Lemma 5.4.
It holds that $V^{k}_{\ell}=\Im(\pi_{\ell})$.
Proof.
Let $\phi=\pi_{\ell}\phi\in\Im(\pi_{\ell})$, and let $\psi\in V^{k}_{0}=\ker B_{\delta}^{k}$ and $\mu\in\Lambda^{\ell}$, $\ell\not=m$. We have
$\widetilde{a}(\phi,\psi)=\kappa_{k}^{-1}\int_{\Gamma_{\ell}^{k}}\lambda(\phi)\psi=0,\qquad b(\phi,\psi)=b(\pi_{\ell}\phi,\psi)=0,$
which implies that $\Im(\pi_{\ell})\subseteq V^{k}_{\ell}$. Let us now prove the converse inequality. Let $\phi\in V_{\ell}^{k}$ and assume that $\pi_{\ell}\phi\not=\phi$. Then $\psi=\phi-\pi_{\ell}\phi\in V_{\ell}^{k}$. We have
$\widetilde{a}(\psi,\xi)=0,\quad\forall\xi\in V^{k}_{0},\qquad\text{ and }b(\psi,\mu)=0,\quad\forall\mu\in\Lambda_{\delta}^{\ell},\ \ell\not=m.$
Together with $b(\psi,\mu)=0$ for all $\mu\in\Lambda_{\delta}^{m}$ this implies $\psi=0$, and therefore $\phi=\pi_{\ell}\phi$, which gives a contradiction.
We can then define the pseudo inverse $B_{\delta}^{+}$ as the solution to (4.16)–(4.18), or, by the reasonong in Remark 4.10, by solving for each $\ell$ the local coupled system (4.21)–(4.23) that takes here the form: find $v_{\ell}^{+}\in\widetilde{V}^{{k^{+}_{\ell}}}_{\ell}$, $v_{\ell}^{-}\in\widetilde{V}^{{k^{-}_{\ell}}}_{\ell}$ such that
(5.9) $\kappa_{k_{\ell}^{+}}\widetilde{a}^{k^{+}_{\ell}}_{\ell}(\widehat{v}_{\ell}^{+},w^{+})-\int_{\Gamma^{k^{+}_{\ell}}_{\ell}}w^{+}\xi_{\ell}=0,\qquad\forall w^{+}\in\widetilde{V}^{k^{+}_{\ell}}_{\ell},$
(5.10) $\kappa_{k_{\ell}^{-}}\widetilde{a}^{k^{-}_{\ell}}_{\ell}(\widehat{v}_{\ell}^{-},w^{-})+\int_{\Gamma^{k^{-}_{\ell}}_{\ell}}w^{-}\xi_{\ell}=0,\qquad\forall v^{-}\in\widetilde{V}^{k^{-}_{\ell}}_{\ell},$
(5.11) $\int_{\Gamma_{\ell}}v^{+}\mu-\int_{\Gamma_{\ell}}v^{-}\mu=\langle\phi,\mu\rangle,\qquad\forall\mu\in\Lambda_{\ell}.$
It is not difficult to ascertain that the solution of (5.9)–(5.11) coincides with the solution to (4.16)–(4.18). We can then solve the former system instead of the latter one, so than we do not need to explicitly compute the spaces $V^{k^{+}_{\ell}}_{\ell}$ and $V^{k^{-}_{\ell}}_{\ell}$.
Thanks to the way we defined $\widetilde{a}$ we immediately see that Assumption 4.7 holds with
$c^{\flat}\simeq C^{\flat}\simeq 1.$
The scalar product $d$ is then defined by (4.15). Moreover, by Lemma 5.3 we have that (4.13) holds with
$\kappa_{*}\simeq\frac{1}{|\log(H/\delta)|^{2}},\qquad K^{*}\simeq 1.$
Following the steps of section 4.1, it follows that the choices above with $\sigma$ defined in (5.4), $\widetilde{a}^{k}$ defined by (5.7) and $V^{k}_{\delta}$ defined by (5.5), or (5.6), leads to a preconditoner $\widehat{M}$ for which Corollary 4.12 gives the bound
$\kappa(\widehat{M}^{-1}\widehat{S}_{h})\lesssim|\log(H/\delta)|^{2}.$
We point out that this preconditioner is, to our knowledge, new. It can be regarded as a dual version of the substructuring preconditioner of [16].
5.2. Domain decomposition with Dirichlet coupling
A second possibility for splitting Problem 5.1 into coupled local problems is to introduce, as an independent variable acting as the multiplier, the trace of the solution on the skeleton $\Gamma$ of the decomposition. The coupled problem then reads: for all $k\in\{1,\ldots,K\}$ find $(u^{k},\zeta^{k})_{k}\in H^{1}(\Omega^{k})\times H^{-1/2}(\Gamma^{k})$ and $\lambda\in H^{1}_{0}(\Omega)|_{\Sigma}$ such that for all $(v^{k},\psi^{k})\in H^{1}(\Omega^{k})\times H^{-1/2}(\Gamma^{k})$
(5.12) $\Big(\int_{\Omega^{k}}\kappa\nabla u^{k}\cdot\nabla v^{k}-\int_{\Gamma^{k}}\zeta^{k}v^{k}+\int_{\Gamma^{k}}u^{k}\psi^{k}\Big)-\int_{\Gamma^{k}}\lambda\psi^{k}=\int_{\Omega^{k}}fv^{k},$
and for all $\mu\in\Lambda$
(5.13) $\sum_{k}\int_{\Gamma^{k}}\zeta^{k}\mu=0.$
Such a problem does not directly fall in our abstract framework, as the local problems (5.12) are not coercive. However, in view of Remark 3.13, we can apply our analysis and solution strategy to a suitably modified, coercive, version of such a problem. As the average $\fint_{\Gamma^{k}}\zeta^{k}$ can be computed directly from the data thanks to the identity
$\int_{\Gamma^{k}}\zeta^{k}=\int_{\Gamma^{k}}\kappa\nabla u^{k}\cdot\nu^{k}=\int_{\Omega^{k}}\nabla\cdot\kappa\nabla u^{k}=-\int_{\Omega^{k}}f,$
we can substitute the unknown $\zeta^{k}$ for its average free component $\phi^{k}=\zeta^{k}-\fint_{\Gamma^{k}}\zeta^{k}$. We then introduce the spaces
$H^{1}_{\text{av}}(\Omega^{k})=\{v\in H^{1}(\Omega^{k}),\ \fint_{\Gamma^{k}}v=0\},\qquad H^{-1/2}_{\text{av}}(\Gamma^{k})=\{\phi\in H^{-1/2}(\Gamma^{k}),\ \langle\phi,1\rangle=0\},$
respectively endowed with the norms
$\|v\|_{H^{1}_{\text{av}}(\Omega^{k})}=\sqrt{\kappa_{k}}|v|_{1,\Omega^{k}},\quad\|\phi\|_{H^{-1/2}_{\text{av}}(\Gamma^{k})}=\frac{1}{\sqrt{\kappa_{k}}}|\phi|_{-1/2,\Gamma^{k}}=\sup_{v\in H^{1}_{av}(\Omega^{k})}\frac{\int_{\Gamma^{k}}\phi v}{|v|_{1,\Omega^{k}}}.$
Then we can rewrite our problem as: look for $(u^{k},\phi^{k})\in V_{k}=H^{1}(\Omega^{k})\times H^{-1/2}_{\text{av}}(\Gamma^{k})$ and $\lambda\in\Lambda=H^{1}(\Omega)|_{\Sigma}$, solution of a problem of the form: for all $(v^{k},\psi^{k})\in V_{k}$, $\mu\in\Lambda$
(5.14) $a(u^{k},\phi^{k};v^{k},\psi^{k})-b(\psi^{k};\lambda)$ $=$ $\langle F;v^{k},\psi^{k}\rangle$
(5.15) $\sum_{k}b(\phi^{k};\mu)$ $=$ $\langle G^{k},\mu\rangle$